OpenStudy (shadowreaper810):

How many H+ ions are present in 10mL of a solution having pH = 10?

OpenStudy (tatianagomezb):

Think of the pH as \[-\log \left[ H ^{+} \right] \] So if the pH= 10 \[\rightarrow - \log \left[ H ^{+} \right]=10\] Then since [H+] it's in Molarity, the result will be in molarity, and it will give you how many moles you have per liter, then you can make the conversion from liters to mLs.

OpenStudy (shadowreaper810):

10^-8 ??

OpenStudy (tatianagomezb):

Once you know the concentration of ions in a pH of ten, that concentration has to be for 1 liter of you solution, remember that concentration is always in mole per liter, then calculate how many of those moles are going to be in 10 mililiters. It's going to be smaller than whatever was the result you got in the initial calculation.

rvc (rvc):

remember the bas is 10

rvc (rvc):

base*

OpenStudy (anonymous):

HOW MANY H+ ions in 10mL of pH = 10. 10^(-10) mol/L = [H+] N = 10^(-10) * 10(-2) N = 10^ (-12) mols 1 mole = 6.02 x 10^23 particles Number of H+= 10^(-12) x (6.02 x 10^23) Number of H+ = 602, 000,000,000 particles of H+